OpenStudy (anonymous):

write the equation of a parallel to 5x-4y=-12, which passes through (8,-1) show steps plz!

OpenStudy (anonymous):

do u know how to do this?

OpenStudy (anonymous):

5x-4y=-12, 5x-4y divided by 5x -12 divided by 5x -4y=-12+5x -4y/-4 -12+5x/-4 y=3+5/-4x parallel -1=8+5/-4x glad i could help and that is correct

OpenStudy (anonymous):

ok awesome thanks

OpenStudy (anonymous):

ur welcome:)